A function is defined on R such that f(x) = (C^2)x when x≀1 and 5Cx-6 when x>1 . Determine the values of C so that f becomes continues on R

1
Expert's answer
2022-04-18T01:46:51-0400

If f(x) is continuous at x=1 then

"\displaystyle\lim_{x\to1}f(x)=f(1)"; "\displaystyle\lim_{x\to1^-}f(x)=f(1)"; "\displaystyle\lim_{x\to1^+}f(x)=f(1)".


"f(x) = \begin{cases}
 C^2x &\text{if } x\le1 \\
 5Cx-6 &\text{if } x>1
\end{cases}"

"\displaystyle\lim_{x\to1^-}f(x)=\displaystyle\lim_{x\to1}C^2x=C^2"

"\displaystyle\lim_{x\to1^+}f(x)=\displaystyle\lim_{x\to1}(5Cx-6)=5C-6"

To make "f(x)" continuous at "x=1" we should find "C" for which

"C^2=5C-6"

"C^2-5C+6=0"

"C=2" ; "C=3" .

Answer: "C=2" or "C=3".